# Hallucinating 360°: Panoramic Street-View Generation via Local Scenes Diffusion and Probabilistic Prompting

### v2

@@ Abstract
Panoramic perception holds significant potential for autonomous driving, enabling vehicles to acquire a comprehensive 360° surround view in a single shot. However, autonomous driving is a data-driven task. Complete panoramic data acquisition requires complex sampling systems and annotation pipelines, which are time-consuming and labor-intensive. Although existing street view generation models have demonstrated strong data regeneration capabilities, they can only learn from the fixed data distribution of existing datasets and cannot achieve high-quality, controllable panoramic generation. In this paper, we propose the first panoramic generation method Percep360 for autonomous driving. Percep360 enables coherent generation of panoramic data with control signals based on the stitched panoramic data. Percep360 focuses on two key aspects: coherence and controllability. Specifically, to overcome the inherent information loss caused by the pinhole sampling process, we propose the Local Scenes Diffusion Method (LSDM). LSDM reformulates the panorama generation as a spatially continuous diffusion process, bridging the gaps between different data distributions. Additionally, to achieve the controllable generation of panoramic images, we propose a Probabilistic Prompting Method (PPM). PPM dynamically selects the most relevant control cues, enabling controllable panoramic image generation. We evaluate the effectiveness of the generated images from three perspectives: image quality assessment (i.e., no-reference and with reference), controllability, and their utility in real-world Bird’s Eye View (BEV) segmentation. Notably, the generated data consistently outperforms the original stitched images in no-reference quality metrics and enhances downstream perception models, leading to an improvement of $2.5\%$ in mIoU for panoramic BEV segmentation. The source code will be publicly available at https://github.com/Bryant-Teng/Percep360.
Index Terms:
Autonomous driving, data generation, world models, panoramic images.
@@ I Introduction
Autonomous vehicles require the integration of surrounding environmental information to achieve 360-degree perception. In this context, panoramic cameras have garnered increasing attention from the research community due to their large Field of View (FoV) [1, 2, 3]. Numerous studies have employed panoramic cameras for downstream perception tasks for street views, such as optical flow estimation [4, 5], object detection [6, 7], and semantic segmentation [8, 9, 1]. However, as a typical data-driven task, developing safe and effective autonomous driving systems heavily relies on the scale and diversity of the collected datasets. Due to the challenges in constructing panoramic data acquisition and annotation systems, the development of panoramic autonomous driving has been hindered by the scarcity of large-scale datasets. OneBEV [10] has attempted to use stitching-based methods to project pinhole images onto a spherical surface, thus simulating panoramic distortion, avoiding the need for data acquisition and offering a new feasible paradigm for panoramic autonomous driving. However, the stitching methods are unable to produce an infinite number of images from prompt input and fail to address the information misalignment caused by the pinhole sampling process.
The recent success of generative models for pinhole cameras has introduced an economic paradigm to avoid the high costs associated with data collection and annotation. Several models [11, 12, 13, 14] have proposed generation frameworks trained on prompt-image pairs, enabling fine-grained control over the diffusion process. In a similar vein, several works [15, 16, 17, 18] have emerged, offering cost-effective solutions for closed-loop autonomous driving data generation. However, existing methods that utilize pinhole samples to facilitate scene and style transfer in driving scenarios fail to ensure coherence in panoramic image synthesis. We remodify the existing models to the panoramic domain. Those networks learn the distribution of stitched images, including misalignment caused by the stitching process and inherent information loss introduced by the pinhole sampling process, as shown in Fig. 1. This can lead to perception errors and localization inaccuracies, further impairing the accuracy of environmental modeling and the reliability of core perception tasks such as object detection and semantic segmentation.
[…]
@@
  • This paper proposes Percep360, the first panoramic generation framework for autonomous driving. Percep360 enables the generation of coherent and controllable panoramic scenarios using inconsistent stitched imagery.
  • To fulfill the inherent information loss caused by the pinhole sampling process, we introduce the Local Scenes Diffusion Method (LSDM). LSDM leverages local spatial information to realize a spatially continuous diffusion process, generating coherent panoramic street scenes.
  • To mitigate the controllability drop that arises when compensating for sample errors, we propose a Probabilistic-Prompt Method (PPM). PPM adjusts the relationships among visual features, textual prompts, and other prompts, resulting in controllable panoramic generation.
@@ II Related Work
Figure 2: The overall architecture of Percep360 leverages the BEV map, textual prompt, depth map, and mask map jointly as guidance signals. With the integration of the LSMD and PMM modules, the framework achieves coherent and controllable panoramic generation.
@@
Panoramic Image Synthesis. The synthesis of panoramic images can be categorized into two aspects: text-based generation and completion based on partial image information. Generation: Text-to-image synthesis [28, 29, 30, 31] has rapidly emerged as a prominent area of research. For instance, PanFusion [32] employs a dual-branch architecture to seamlessly integrate additional constraints, enabling more customizable and high-quality panoramic outputs. Several works [33, 34, 35] achieve panoramic image completion by extending the field of view, enabling seamless and realistic scene expansion. Completion: The image inpainting task aims to restore missing regions of an image while maintaining overall coherence. PICFormer [36] introduces the Codebook-based sharing paradigm and enhances restoration performance through plus-realism image completion [37]. RePaint [38] models the semantics and texture of missing regions by leveraging a denoising diffusion probabilistic model. BrushNet [39, 40] incorporates a mask branch to better distinguish masked regions. Several works [41, 42] introduce the rotation-based paradigm to address the continuity challenges in panoramic image completion, ensuring smoother and more coherent results. However, a fundamental limitation exists: These works are trained using consistent images paired with text descriptions. Such a paradigm is not suitable for generating coherent panoramic data from stitched images. Moreover, existing panoramic image synthesis methods utilize weak textual prompts, which fail to meet the precise requirements of autonomous driving for accurate road topology representation.
@@ III Methodology
Figure 3: The architecture of LSMD and PPM is illustrated in Figures (a) and (b), respectively. By applying circular rotations to diverse prompts and BEV features, LSDM reformulates panoramic image generation as a spatially continuous diffusion process. PPM further enhances controllability by dynamically selecting the most relevant control cues, leading to accurate and consistent panoramic image synthesis.
@@ III-A Preliminary and Problem Formulation
Preliminary: Latent Diffusion Models (LDM) [43] represent traditional diffusion models by conducting the diffusion process within a compact latent representation space, typically involving a two-step procedure: feature compression and iterative denoising for modeling data distributions. In the feature compression, input images in the pixel space are first transformed into compact latent embeddings using a perceptual compression model $E$, significantly reducing the complexity associated with processing high-dimensional data: $z=E(x)$. This latent representation $z$ passes through the diffusion process and can be decoded back into the original image domain using the decoder $D$, formally expressed as: $\hat{x}={D}(z)$. In the iterative denoising for modeling data distributions, the diffusion process is carried out within this latent space. In the forward diffusion stage, gaussian noise is gradually added to the latent embeddings, progressively transforming the latent representation into a distribution close to a standard Gaussian: $z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\enspace\epsilon\sim\mathcal{N}(0,I)$. Here, $\bar{\alpha}_{t}$ is a pre-defined noise scheduling parameter, and $t$ denotes the diffusion timestep. In the reverse denoising stage, a U-Net is trained to estimate and remove the added noise from the noisy latent representation $z_{t}$. This network is parameterized as $\epsilon_{\theta}$, trained by minimizing the following mean-square error loss: $\min_{\theta}\mathbb{E}_{z,c,\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|^{2}\right]$, where $c$ represents conditional information, guiding the conditional image generation. After training, the LDM can generate samples by iterative denoising a latent feature initialized from Gaussian noise. Formally, this iterative sampling process can be expressed as: $p_{\theta}(z_{t-1}|z_{t},c)=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,c),\Sigma_{\theta}(z_{t},t,c))$, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are predicted by the denoising network.
@@
Problem Formulation: Given a panoramic image defined over a pixel domain $\Omega\subset\mathbb{R}^{2}$, the entire data distribution of the image is denoted by $f(x)$, where $x\in\Omega$. The panoramic image is constructed by stitching multiple pinhole camera views, resulting in a composite pixel space partitioned into twelve regions: six coherent regions $\{\mathcal{R}_{i}\}_{i=1}^{6}$ and six aliased regions $\{\mathcal{A}_{i}\}_{i=1}^{6}$. Thus, the domain can be expressed as: $\Omega=\left(\bigcup_{i=1}^{6}\mathcal{R}_{i}\right)\cup\left(\bigcup_{i=1}^{6}\mathcal{A}_{i}\right),\enspace\mathcal{R}_{i}\cap\mathcal{A}_{j}=\emptyset,\enspace\forall i,j$. In the initial state, the distribution of the image exhibits regional coherence. The local distribution over each coherent region $\mathcal{R}_{i}$, denoted $f_{i}(x)$, possesses strong spatial consistency. In contrast, the distribution over each aliased region $\mathcal{A}_{i}$, denoted $\tilde{f}_{i}(x)$, is complex and discontinuous due to stitching artifacts. Our objective is to enhance the model’s ability to capture the structure of the coherent regions by introducing an operation $R_{\theta}(\cdot)$. Meanwhile, by incorporating a control method, the network leverages prompt signals $A$ and $B$ to generate a controllable distribution: $f(x\mid A,B)\approx\sum_{i=1}^{6}\mathbb{I}_{\mathcal{R}_{i}}(x)\cdot f_{i}(x\mid A,B)$, where $\mathbb{I}_{\mathcal{R}_{i}}(x)$ is the indicator function for region $\mathcal{R}_{i}$. Specifically, we seek to transform the initial discontinuous mixed distribution: $f(x)=\sum_{i=1}^{6}\mathbb{I}_{\mathcal{R}_{i}}(x)\cdot f_{i}(x)+\sum_{i=1}^{6}\mathbb{I}_{\mathcal{A}_{i}}(x)\cdot\tilde{f}_{i}(x)$, into a coherent, structure-dominant distribution: $f(x\mid A,B)\longrightarrow\tilde{f}_{\text{coherent}}(x),\enspace x\in\Omega$.
@@ III-B The Structure of Percep360
In this paper, we propose Percep360, the first paradigm for achieving high-quality controllable panoramic image generation. Specifically, Percep360 adopts a side-controlling architecture, following [11, 12]. The network introduces LSMD, enabling more coherent image generation under the panorama stitching setting. Furthermore, we introduce the PPM module, which injects prompt embeddings into the global features of panoramic images to enable controllable image generation.
@@
Next, the BEV map is encoded using a dedicated BEV encoder $\mathcal{E}_{\text{BEV}}$, following BEVControl [14]. The resulting feature map is resized to match the spatial resolution of the 2D features:
@@
The design of LSMD: To achieve more coherent panoramic generation, we propose the Local Scenes Diffusion Method (LSMD), as shown in Fig. 3, which mitigates the discontinuities at image seams. Since a panoramic image spans a full $360^{\circ}$ horizontal field of view, we treat its domain as a spherical manifold where the left and right boundaries are adjacent. For the 2D inputs, including the image $\mathbf{I}$, mask $\mathbf{M}$, and depth map $\mathbf{D}$, we perform a horizontal wrap-around rotation before feeding them into the network. A rotation angle $\theta\in[0,2\pi)$ is randomly sampled and converted into a pixel displacement:
@@
For the BEV input $\mathbf{B}$, we encode it to obtain the feature map:
@@
This feature-space rotation is consistent with the 2D image rotation, but applied directly to the encoded BEV representation. By synchronizing the angular shift across both 2D and BEV modalities, LSMD facilitates geometrically aligned and spatially consistent panoramic generation.
@@
NRIQA IQA Control Rank.
Oracle 19.39 12.06 0.59 0.47 N/A
Percep360 20.24 (1) 11.44 (1) 0.16 (1) 14.43 (2) 0.25 (1) 0.13 (1) 7 (1)
Table I: Comparison with state-of-the-art methods on the nuScenes validation set in Generation Quality. $\uparrow$/$\downarrow$ / ↓ indicates that a higher/lower value is better. ‘Oracle’ presents the row results of corresponding models trained on the nuScenes training set. indicates that the model has undergone structural modifications to accommodate the panoramic image generation task.
Figure 4: Visualizations of the generation results for the ground truth, BEVControl [14], and MagicDrive [18] are provided. Regions of stitching misalignment are highlighted with yellow boxes. It can be observed that, due to the lack of coherent generation capability, existing methods produce discontinuous results. In contrast, our method achieves significantly improved coherence without compromising controllability.
@@
Benefiting from the proposed design, our method enhances the quality of generated images and has the potential to boost the performance downstream panoramic BEV perception.

### v3

@@ Abstract
Panoramic perception holds significant potential for autonomous driving, enabling vehicles to acquire a comprehensive 360° surround view in a single shot. However, autonomous driving is a data-driven task. Complete panoramic data acquisition requires complex sampling systems and annotation pipelines, which are time-consuming and labor-intensive. Although existing street view generation models have demonstrated strong data regeneration capabilities, they can only learn from the fixed data distribution of existing datasets and cannot leverage stitched pinhole images as a supervisory signal. In this paper, we propose the first panoramic generation method Percep360 for autonomous driving. Percep360 enables coherent generation of panoramic data with control signals based on the stitched panoramic data. Percep360 focuses on two key aspects: coherence and controllability. Specifically, to overcome the inherent information loss caused by the pinhole sampling process, we propose the Local Scenes Diffusion Method (LSDM). LSDM reformulates the panorama generation as a spatially continuous diffusion process, bridging the gaps between different data distributions. Additionally, to achieve the controllable generation of panoramic images, we propose a Probabilistic Prompting Method (PPM). PPM dynamically selects the most relevant control cues, enabling controllable panoramic image generation. We evaluate the effectiveness of the generated images from three perspectives: image quality assessment (i.e., no-reference and with reference), controllability, and their utility in real-world Bird’s Eye View (BEV) segmentation. Notably, the generated data consistently outperforms the original stitched images in no-reference quality metrics and enhances downstream perception models, leading to an improvement of $2.5\%$ in mIoU for panoramic BEV segmentation. The source code will be publicly available at https://github.com/FeiT-FeiTeng/Percep360.
@@ I Introduction
Autonomous vehicles require the integration of surrounding environmental information to achieve 360-degree perception. In this context, panoramic cameras have garnered increasing attention from the research community due to their large Field of View (FoV) [7, 42, 1]. Numerous studies have employed panoramic cameras for downstream perception tasks for street views, such as optical flow estimation [27, 15], object detection [41, 19], and semantic segmentation [7, 50, 52]. However, as a typical data-driven task, developing safe and effective autonomous driving systems heavily relies on the scale and diversity of the collected datasets. Due to the challenges in constructing panoramic data acquisition and annotation systems, the development of panoramic autonomous driving has been hindered by the scarcity of large-scale datasets. OneBEV [35] has attempted to use stitching-based methods to project pinhole images onto a spherical surface, thus simulating panoramic distortion, avoiding the need for data acquisition and offering a new feasible paradigm for panoramic autonomous driving. However, the stitching methods are unable to produce an infinite number of images from prompt input and fail to address the information misalignment caused by the sampling process.
Figure 1: Percep360 leverages diverse prompts to generate coherent panoramic images. The generated images exhibit improved quality and controllability. As a data generator, Percep360 synthesizes diverse data to avoid the expensive process of data processing, thereby enhancing the performance of semantic mapping.
The recent success of generative models for pinhole cameras has introduced an economic paradigm to avoid the high costs associated with data collection and annotation. Several models [47, 23, 29, 44] have proposed generation frameworks trained on prompt-image pairs, enabling fine-grained control over the diffusion process. In a similar vein, several works [33, 51, 36, 6] have emerged, offering cost-effective solutions for closed-loop autonomous driving data generation. However, existing methods that utilize pinhole samples to facilitate scene and style transfer in driving scenarios fail to ensure coherence in panoramic image synthesis. We remodify the existing models for the panoramic domain. Those networks learn the distribution of stitched images, including misalignment caused by the stitching process and inherent information loss introduced by the pinhole sampling process, as shown in Fig. 2. This can lead to perception errors and localization inaccuracies, further impairing the accuracy of environmental modeling and ultimately affecting the practical applications of robots.
Thus, effectively generating coherent panoramic images remains a crucial challenge for panoramic autonomous driving systems.
@@
  • This paper proposes Percep360, the first panoramic generation framework for autonomous driving. Percep360 enables the generation of coherent and controllable panoramic data using inconsistent stitched imagery.
  • To fulfill the inherent information loss caused by the pinhole sampling process, we introduce the Local Scenes Diffusion Method (LSDM). LSDM leverages local information to realize a spatially continuous diffusion process, generating coherent panoramic scenes.
  • To mitigate the controllability drop that arises when compensating for sample errors, we propose a Probabilistic-Prompt Method (PPM). PPM adjusts the relationships among visual, textual, and other prompts, resulting in controllable panoramic generation.
@@ II Related Work
Figure 3: The overall architecture of Percep360 leverages the BEV map, textual prompt, depth map, and mask map jointly as guidance signals. With the integration of the LSDM and PPM modules, the framework achieves coherent and controllable panoramic generation.
Panoramic Image Synthesis. The synthesis of panoramic images can be categorized into two aspects: text-based generation and completion based on partial image information. Generation: Text-to-image synthesis [22, 2, 9, 43] has rapidly emerged as a prominent area of research. For instance, PanFusion [45] employs a dual-branch architecture to seamlessly integrate additional constraints, enabling more customizable and high-quality panoramic outputs. Several works [16, 12, 32] achieve panoramic image completion by extending the field of view, enabling seamless and realistic scene expansion. Completion: The image inpainting task aims to restore missing regions of an image while maintaining overall coherence. PICFormer [49] introduces the Codebook-based sharing paradigm and enhances restoration performance through plus-realism image completion [48]. RePaint [18] models the semantics and texture of missing regions by leveraging a denoising diffusion probabilistic model. BrushNet [11, 14] incorporates a mask branch to better distinguish masked regions. Several works [30, 38] introduce the rotation-based paradigm to address the continuity challenges in panoramic image completion, ensuring smoother and more coherent results. However, a fundamental limitation exists: These works are trained using consistent images. Such a paradigm is not suitable for generating coherent panoramic data from imperfect samples. Moreover, existing panoramic image synthesis methods utilize weak textual prompts, which fail to meet the requirements of autonomous driving for accurate road topology representation.
@@ III-A Preliminary and Problem Formulation
Preliminary: Latent Diffusion Models (LDM) [25] represent traditional diffusion models by conducting the diffusion process within a compact latent representation space, typically involving a two-step procedure: feature compression and iterative denoising for modeling data distributions. In the feature compression, input images in the pixel space are first transformed into compact latent embeddings using a perceptual compression model $E$, significantly reducing the complexity associated with processing high-dimensional data: $z=E(x)$. This latent representation $z$ passes through the diffusion process and can be decoded back into the original image domain using the decoder $D$, formally expressed as: $\hat{x}={D}(z)$. In the iterative denoising for modeling data distributions, the diffusion process is carried out within this latent space. In the forward diffusion stage, gaussian noise is gradually added to the latent embeddings, progressively transforming the latent representation into a distribution close to a standard Gaussian: $z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\kern 5.0pt\epsilon\sim\mathcal{N}(0,I)$. Here, $\bar{\alpha}_{t}$ is a pre-defined noise scheduling parameter, and $t$ denotes the diffusion timestep. In the reverse denoising stage, a U-Net is trained to estimate and remove the added noise from the noisy latent representation $z_{t}$. This network is parameterized as $\epsilon_{\theta}$, trained by minimizing the following mean-square error loss: $\min_{\theta}\mathbb{E}_{z,c,\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|^{2}\right]$, where $c$ represents conditional information, guiding the conditional image generation. After training, the LDM can generate samples by iterative denoising a latent feature initialized from Gaussian noise. Formally, this iterative sampling process can be expressed as: $p_{\theta}(z_{t-1}|z_{t},c)=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,c),\Sigma_{\theta}(z_{t},t,c))$, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are predicted by the denoising network.
Figure 4: The architecture of LSDM and PPM is illustrated in Figures (a) and (b), respectively. By applying circular rotations to diverse prompts and BEV features, LSDM reformulates panoramic image generation as a spatially continuous diffusion process. PPM further enhances controllability by dynamically selecting the most relevant control cues, leading to an accurate and consistent panoramic image.
@@
Problem Formulation: Given a panoramic image defined over a pixel domain $\Omega\subset\mathbb{R}^{2}$, the entire data distribution of the image is denoted by $f(x)$, where $x\in\Omega$. The panoramic image is constructed by stitching multiple pinhole camera views, resulting in a composite pixel space partitioned into twelve regions: six coherent regions $\{\mathcal{R}_{i}\}_{i=1}^{6}$ and six aliased regions $\{\mathcal{A}_{i}\}_{i=1}^{6}$. Thus, the domain can be expressed as: $\Omega=\left(\bigcup_{i=1}^{6}\mathcal{R}_{i}\right)\cup\left(\bigcup_{i=1}^{6}\mathcal{A}_{i}\right),\kern 5.0pt\mathcal{R}_{i}\cap\mathcal{A}_{j}=\emptyset,\kern 5.0pt\forall i,j$. In the initial state, the distribution of the image exhibits regional coherence. The local distribution over each coherent region $\mathcal{R}_{i}$, denoted $f_{i}(x)$, possesses strong spatial consistency. In contrast, the distribution over each aliased region $\mathcal{A}_{i}$, denoted $\tilde{f}_{i}(x)$, is complex and discontinuous due to stitching artifacts. Our objective is to enhance the model’s ability to capture the structure of the coherent regions by introducing an operation $R_{\theta}(\cdot)$. Meanwhile, by incorporating a control method, the network leverages prompt signals $A$ and $B$ to generate a controllable distribution: $f(x\mid A,B)\approx\sum_{i=1}^{6}\mathbb{I}_{\mathcal{R}_{i}}(x)\cdot f_{i}(x\mid A,B)$, where $\mathbb{I}_{\mathcal{R}_{i}}(x)$ is the indicator function for region $\mathcal{R}_{i}$. Specifically, we seek to transform the initial discontinuous mixed distribution: $f(x)=\sum_{i=1}^{6}\mathbb{I}_{\mathcal{R}_{i}}(x)\cdot f_{i}(x)+\sum_{i=1}^{6}\mathbb{I}_{\mathcal{A}_{i}}(x)\cdot\tilde{f}_{i}(x)$, into a coherent, structure-dominant distribution: $f(x\mid A,B)\longrightarrow\tilde{f}_{\text{coherent}}(x),\kern 5.0ptx\in\Omega$.
@@ III-B The Structure of Percep360
In this paper, we propose Percep360, the first paradigm for achieving high-quality controllable panoramic image generation. Specifically, Percep360 adopts a side-controlling architecture, following [47, 23]. The network introduces LSDM, enabling more coherent image generation under the panorama stitching setting. Furthermore, we introduce the PPM module, which injects prompt embeddings into the global features of panoramic images to enable controllable image generation.
@@
Next, the BEV map is encoded using a dedicated BEV encoder $\mathcal{E}_{\text{BEV}}$, following BEVControl [44]. The resulting feature map is resized to match the spatial resolution of the 2D features, as shown in the following equation:
@@
The design of LSDM: To achieve more coherent panoramic generation, we propose the Local Scenes Diffusion Method (LSDM), as shown in Fig. 4, which mitigates the discontinuities at image seams. Since a panoramic image spans a full $360^{\circ}$ horizontal field of view, we treat its domain as a spherical manifold where the left and right boundaries are adjacent. For the 2D inputs, including the image $\mathbf{I}$, mask $\mathbf{M}$, and depth map $\mathbf{D}$, we perform a horizontal wrap-around rotation before feeding them into the network. A rotation angle $\theta\in[0,2\pi)$ is randomly sampled and converted into a pixel displacement:
@@
The BEV input $\mathbf{B}$ is encoded into a feature map:
@@
This feature-space rotation is consistent with the 2D image rotation, but applied directly to the encoded BEV representation. By synchronizing the angular shift across both 2D and BEV modalities, LSDM facilitates geometrically aligned and spatially consistent panoramic generation.
NRIQA IQA Control Rank.
[…]
Percep360 20.24 (1) 11.44 (1) 0.16 (1) 14.43 (2) 0.25 (1) 0.13 (1) 7 (1)
Table I: Comparison with state-of-the-art methods on the nuScenes validation set in Generation Quality. $\uparrow/\downarrow$ indicates that a higher/lower value is better. ‘Oracle’ presents the row results of corresponding models trained on the nuScenes training set. indicates that the model has undergone structural modifications to accommodate the panoramic image generation task.
@@
Figure 5: Visualizations of the generation results for the ground truth, BEVControl [44], and MagicDrive [6] are provided. Regions of stitching misalignment are highlighted with yellow boxes. It can be observed that, due to the lack of coherent generation capability, existing methods produce discontinuous results. In contrast, our method achieves improved coherence without compromising controllability.
@@
Benefiting from the proposed design, our method enhances the quality of generated images and has the potential to boost the performance of downstream panoramic BEV perception.